I am so confused with this problem. Can someone show me how this is done?
Find values for m and b in the following system so that the solution to the system is (-3, 4).
5x + 7y = b
mx+ y = 22

Answer by ptaylor(1894)

5x + 7y = b--------------eq1
mx+ y = 22----------------eq2
Since we want the solution to this system of equations to be (-3,4), we substitute (-3,4) into each equation and solve for the unknowns.
From eq1
5*(-3)+7*4=b
b=-15+28
b=13

From eq2
-3m+4=22
-3m=22-4
-3m=18
m=-6
So---

5x+7y=13-------------------Final eq1
-6x+y=22-------------------------Final eq2
CK
Substitute (-3,4) into eq 1
-15+28=13
13=13------ok
Substitute (-3,4) into eq2
-6(-3)+4=22
18+4=22
22=22-------ok

Hope this helps---ptaylor